20) You are required to dilute a stock solution of 18.1 M H¬2SO4 to 300 mL of 3.0 M H2SO4. How many mL of stock are required? How many mL of water?
Answer #1

molarity of stock solution, M1 = 18.1 M

let the volume of stock solution needed, V1 = V mL

molarity of diluted solution, M2 = 3.0 M

volume of diluted solution, V2 = 300 mL

now, by using the general dilution equation,

M1V1 = M2V2

18.1 * V = 3.0 * 300

or V = (3*300)/18.1 mL = 49.7237 mL

now we have to add water to this volume of stock solution to make it 300 mL.

that is needed volume of water = 300mL-49.7237mL = 250.2763 mL

needed volume of stock solution = 49.72 mL

needed volume of water = 250.28 mL